OpenStudy (anonymous):

7r-5s=44 5r+7s=42 I need help finding the solution please?

OpenStudy (anonymous):

Can you please help me work it out so that I can understand please.

OpenStudy (anonymous):

(7r - 5s =44)*7 to 49r -35s = 308 (5r +7s = 42)*5 to 25r +35s = 210 Therefore: 74r = 518 r =7 & s = 1

OpenStudy (anonymous):

so the order pair would be (308,210) in that order.

OpenStudy (anonymous):

no, the order pair would be (7,1) for (r,s)

OpenStudy (anonymous):

o ok I see where you got that I was confused now I am not.

OpenStudy (anonymous):

I also have y+x=-9 I am trying to get y by it's self